### v1

# Are Data Augmentation Methods in Named Entity Recognition Applicable for Uncertainty Estimation?

Abstract
This work investigates the impact of data augmentation on confidence calibration and uncertainty estimation in Named Entity Recognition (NER) tasks. For the future advance of NER in safety-critical fields like healthcare and finance, it is essential to achieve accurate predictions with calibrated confidence when applying Deep Neural Networks (DNNs), including Pre-trained Language Models (PLMs), as a real-world application. However, DNNs are prone to miscalibration, which limits their applicability. Moreover, existing methods for calibration and uncertainty estimation are computational expensive. Our investigation in NER found that data augmentation improves calibration and uncertainty in cross-genre and cross-lingual setting, especially in-domain setting. Furthermore, we showed that the calibration for NER tends to be more effective when the perplexity of the sentences generated by data augmentation is lower, and that increasing the size of the augmentation further improves calibration and uncertainty.
Are Data Augmentation Methods in Named Entity Recognition Applicable for Uncertainty Estimation?
Wataru Hashimoto,  Hidetaka Kamigaito,  Taro Watanabe Nara Institute of Science and Technology {hashimoto.wataru.hq3, kamigaito.h, taro}@is.naist.jp
1 Introduction
Named Entity Recognition (NER) is a one of the fundamental tasks in Natural Language Processing (NLP) to find mentions of named entities and classify them into predefined categories. The predicted information by NER is essential for downstream tasks like event detection Vavliakis et al. (2013), information retrieval Cowan et al. (2015), and masking of personal user information Kodandaram et al. (2021). Due to the demand, NER is the underlying technology for information extraction from text and documents.
Based on the recent advances in Deep Neural Networks (DNNs), NER’s performance is also improved like other NLP fields. In recent years, Pre-trained Language Models (PLMs) based architectures, such as BERT Devlin et al. (2019) and DeBERTa He et al. (2021), have been strong baselines in many NLP tasks, including NER.
[…]
Confidence calibration and uncertainty estimation methods are ways to solve the miscalibration of DNNs, and have been applied in NLP tasks such as text classification Xiao and Wang (2019), structured prediction Jiang et al. (2022); Reich et al. (2020), question answering Si et al. (2022), and machine translation Malinin and Gales (2021). However, many methods for confidence calibration and uncertainty estimation, typically Monte-Carlo Dropout (MC Dropout) Gal and Ghahramani (2016), are computationally expensive due to multiple stochastic inferences, making them difficult for real-world application.
Data augmentation has also been applied for NER Dai and Adel (2020); Zhou et al. (2022), though, it was focusing on the generalization ability on low-resource data. In computer vision (CV) areas, data augmentation makes the model more robust to the input and leads to confidence calibrations Wen et al. (2021); Liu et al. (2023), in which the same labels are trained on different representations of the input than the original data. Based on the findings of these previous studies, there is a possibility that data augmentation in NER can improve confidence calibration without increasing inference time, in contrast to the conventional confidence calibration and uncertainty estimation methods.
In this study, we conducted comprehensive experiments to analyze the impact of data augmentation methods for NER Dai and Adel (2020); Zhou et al. (2022) on the confidence calibration and uncertainty in the cross-genre and cross-lingual settings on OntoNotes 5.0 Pradhan et al. (2013) and MultiCoNER Malmasi et al. (2022), respectively.
Our experiments yield several findings. First, some data augmentation methods in NER lead to improved confidence calibration and uncertainty estimation, especially in-domain. In particular, entity-prediction-based data augmentation Zhou et al. (2022) and entity replacement from the same entity type Dai and Adel (2020) show good performance. On the other hand, common confidence calibration methods, MC Dropout or TS Guo et al. (2017) have worse confidence calibration and uncertainty estimation performance than the data augmentation methods in NER, even though the data augmentation methods do not aim to improve confidence calibration and uncertainty estimation. Moreover, increasing the augmentation size improves performance in confidence calibration and uncertainty estimation. The improvement tends to be better the lower the perplexity of the sentences generated by the data augmentation. Our code will be released after acceptance.
2 Related Work
Named Entity Recognition
In the last decade, NER using DNNs has been widely successful; Lample et al. (2016) reported a sequence-labeling model combining bi-directional LSTM with CRF (BiLSTM-CRF). Akbik et al. (2018) proposed contextualized character-level word embeddings combined with BiLSTM-CRF. In recent years, NER models based on PLMs, such as BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and DeBERTa He et al. (2021), have achieved state-of-the-art performance.
Uncertainty Estimation
In general, DNNs are prone to miscalibration and overconfidence Guo et al. (2017) especially without pretraining Desai and Durrett (2020); Ulmer et al. (2022). One way to estimate uncertainty is to run multiple stochastic predictions. Deep Ensemble Lakshminarayanan et al. (2017) trains multiple DNN models and integrates their multiple stochastic predictions to make a final prediction. MC Dropout Gal and Ghahramani (2016) applies Dropout Srivastava et al. (2014) regularization at both training and inference time, and by taking multiple samples of the network outputs during inference. These are known to perform calibration well in many cases Ovadia et al. (2019); Immer et al. (2021), but their practical use is hampered by the fact that they make multiple probabilistic predictions. A relatively lightweight calibration method is the post-hoc approach. For example, temperature scaling Guo et al. (2017) performs calibration via dividing logits by a constant, which is a simple and lightweight baseline.
Data Augmentation
Data augmentation methods are widely used in machine learning, CV, and NLP areas. More recent attention has focused on the provision of data augmentation methods to improve calibration and uncertainty. Test-time augmentation (TTA) Ashukha et al. (2020) generates multiple samples during inference and integrates the predictions to estimate the prediction uncertainty. MixUp Zhang et al. (2018) uses linear interpolation between two samples to augment a new sample with soft labels, which has been investigated for situations where it is effective for calibration Zhang et al. (2022).
In NLP tasks, the impact of data augmentation on calibration in text classification has been investigated in recent study Kim et al. (2023), but only for In-domain (ID) and not for NER. Furthermore, it has been found that predictive performance is driven by data augmentation in NER Dai and Adel (2020); Chen et al. (2020); Zhou et al. (2022); Chen et al. (2022); Hu et al. (2023), but these studies have focused only on the predictive performance of NER and have not evaluated for calibration and uncertainty. This is the first study to comprehensively investigate the impact of data augmentation on calibration and uncertainty in NER, both in ID and OOD (Out-of-domain) settings.
3 Methods
In this section, we describe the popular baseline methods for confidence calibration and data augmentation methods for NER. Details about existing calibration methods are described in Appendix B.
3.1 Existing Calibration Methods
Baseline
Baseline uses the maximum probability from the softmax layer.
Temperature Scaling (TS)
TS Guo et al. (2017) is a post-processing technique for calibrating the confidence scores outputted by a neural network. It involves scaling the logits (i.e., the outputs of the final layer before the softmax) by a temperature parameter $T$ before applying the softmax function to obtain the calibrated probabilities.
Label Smoothing (LS)
LS Miller et al. (1996); Pereyra et al. (2017) is prevalent regularization technique in machine learning, introduces a controlled level of uncertainty into the training process by modifying the cross-entropy loss.
Monte-Carlo Dropout (MC Dropout)
MC Dropout is a regularization technique that can be used for uncertainty estimation in neural networks, which requires multiple stochastic inferences Gal and Ghahramani (2016). We perform 20 stochastic inferences and output their average.
3.2 Data Augmentation Methods for NER
We investigate data augmentation methods in NER Dai and Adel (2020); Zhou et al. (2022) for confidence calibration and uncertainty estimation.
Label-wise Token Replacement (LwTR)
LwTR uses binomial distribution to determine whether a token is replaced. The chosen token is randomly replaced with another token with the same label based on label-wise token distribution on training data. Thus, LwTR keeps the original label sequence.
Mention Replacement (MR)
Unlike LwTR, MR replaces an entity with another entity with the same label instead of a token. Other parts are the same as LwTR. Since entities can have multiple tokens, MR does not keep the original label sequence.
Synonym Replacement (SR)
SR is similar to LwTR except that SR replaces a token with its synonym in WordNet Miller (1995). Since the synonym can have multiple tokens, SR does not keep the original label sequence.
Masked Entity Language Modeling (MELM)
MELM Zhou et al. (2022) performs data augmentation using a language model that predicts contextually appropriate entities for sentences in which entity parts are masked by entity markers.
4 Evaluation Metrics
We use Expected Calibration Error (ECE), Maximum Calibration Error (MCE), and Area Under Precision-Recall Curve (AUPRC) to evaluate confidence calibration and uncertainty estimation.
4.1 Expected Calibration Error (ECE)
ECE Naeini et al. (2015) measures the difference between the accuracy and confidence of a model. Specifically, it calculates the difference between the average confidence and the actual accuracy of the model on different confidence levels. Formally, ECE is defined as:
$\text{ECE}=\sum_{b=1}^{B}\frac{|\mathcal{D}_{b}|}{n}\left\lvert\text{acc}(\mathcal{D}_{b})-\text{conf}(\mathcal{D}_{b})\right\rvert$
where $B$ is the number of confidence interval bins, $\mathcal{D}_{b}$ is the set of examples whose predicted confidence scores fall in the $b$-th interval, $n$ is the total number of examples, $\text{acc}(\mathcal{D}_{b})$ is the accuracy of the model on the examples in $\mathcal{D}_{b}$, and $\text{conf}(\mathcal{D}_{b})$ is the average confidence of the model on the examples in $\mathcal{D}_{b}$.
4.2 Maximum Calibration Error (MCE)
MCE Naeini et al. (2015) is the maximum difference between the accuracy and the confidence of the model on different confidence levels. Formally, MCE is defined as:
$\text{MCE}=\max_{b=1}^{B}\left\lvert\text{acc}(\mathcal{D}_{b})-\text{conf}(\mathcal{D}_{b})\right\rvert,$
MCE takes the maximum calibration error in each bin, not the expectation; a smaller MCE means that the model’s predictions are less likely to be far off in a given confidence region.
4.3 Area Under the Precision-Recall Curve (AUPRC)
AUPRC is the summary statistic the relationship between precision and recall at different thresholds. The higher the value, the higher the overall precision at a given threshold.
5 Experimental Settings
5.1 Datasets
We conducted experiments on two different NER datasets to evaluate the performance of confidence calibration methods in different settings. For the cross-genre evaluation, we used the OntoNotes 5.0 dataset Pradhan et al. (2013), which consists of six different genres, broadcast conversation ($\mathtt{bc}$), broadcast news ($\mathtt{bn}$), magazine ($\mathtt{mz}$), newswire ($\mathtt{nw}$), telephone conversation ($\mathtt{tc}$), and web data ($\mathtt{wb}$). This dataset is commonly used for NER evaluation in a cross-domain setting Chen et al. (2021).
For the cross-lingual evaluation, we used the MultiCoNER dataset, which is a large multilingual NER dataset from Wikipedia sentences, questions, and search queries Malmasi et al. (2022). We selected English as the source language and English, German, Spanish, Hindi, and Bangla as the target languages. The details of the dataset statistics are provided in Table 1.
Dataset &\&& Domain $N_{ent}$ Train Dev Test
OntoNotes 5.0
$\mathtt{bc}$ 18 11,866 2,117 2,211
$\mathtt{bn}$ 18 10,683 1,295 1,357
$\mathtt{mz}$ 18 6,911 642 780
$\mathtt{nw}$ 18 33,908 5,771 2,197
$\mathtt{tc}$ 18 11,162 1,634 1,366
$\mathtt{wb}$ 18 7,592 1,634 1,366
MultiCoNER
English ($\mathtt{EN}$) 6 15,300 800 10,000
German ($\mathtt{DE}$) 6 - - 10,000
Spanish ($\mathtt{ES}$) 6 - - 10,000
Hindi ($\mathtt{HI}$) 6 - - 10,000
Table 1: Dataset statistics. The table presents the number of entity types, and sequences for the train, development, and test parts of the datasets. For MultiCoNER, we randomly sampled and fixed 10,000 cases out of 200,000 test cases.
5.2 Training Details
In all experiments, we train out models on a single NVIDIA A100 GPU with 40GB of memory. We used MIT-licensed mDeBERTaV3 (microsoft/mdeberta-v3-base) He et al. (2023) whose model size is 278M, as a multilingual transformer encoder from Hugging Face $\mathtt{transformers}$ Wolf et al. (2020) pre-trained model checkpoints, and extracted entities via sequence labeling. Cross-entropy loss is minimized by AdamW Loshchilov and Hutter (2019) with a linear scheduler Goyal et al. (2017). The batch size is 32, and gradient clipping is applied with maximum norm of 1. The initial learning rate was set to 1e-5. To avoid overfitting, we also applied early stopping with $patients=5$.
For the temperature parameter in TS, we used Optuna Akiba et al. (2019) to optimize the temperature parameter based on dev set loss with a search range of [0.001, 0.002, …, 5.000] in 100 trials. In addition, we optimized the binomial distribution parameter for data augmentation methods using the dev set by a grid search in the range of [0.1, 0.2, …, 0.8]. In LS, we conducted a grid search in the range of [0.01, 0.05, 0.1, 0.2, 0.3] to optimize the smoothing parameter. In the case of MELM, mask rate $\eta$ during fine tuning and mask parameter $\mu$ during generation are hyperparameters. We conducted a grid search for each hyperparameter in the range [0.3, 0.5, 0.7], as in Zhou et al. (2022).
We perform each experiment 10 times using different random seeds, collect evaluation metric values, and report their average and standard deviation. For convenience, the reported values are multiplied by 100.
Methods $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓)
Baseline 18.87$\pm$0.73 23.58$\pm$1.01 11.50$\pm$0.75 16.14$\pm$1.97 15.75$\pm$0.94 20.93$\pm$0.97 11.74$\pm$0.27 16.15$\pm$0.77 31.17$\pm$1.56 33.81$\pm$1.67 28.86$\pm$1.51 34.38$\pm$1.82
TS 18.86$\pm$0.68 23.22$\pm$0.86 11.25$\pm$0.55 15.43$\pm$1.41 15.40$\pm$0.74 20.30$\pm$1.23 11.71$\pm$0.36 15.80$\pm$0.85 27.95$\pm$2.51 30.70$\pm$2.55 29.70$\pm$1.54 34.88$\pm$1.66
LS 19.29$\pm$1.04 24.11$\pm$1.57 11.42$\pm$0.52 15.31$\pm$1.24 15.59$\pm$0.85 20.91$\pm$1.30 12.05$\pm$0.20 16.83$\pm$0.36 26.46$\pm$1.36 28.89$\pm$1.42 29.34$\pm$2.25 34.86$\pm$2.22
MCDropout 18.69$\pm$0.71 23.54$\pm$1.31 11.38$\pm$0.71 15.73$\pm$1.60 15.89$\pm$0.29 21.15$\pm$0.54 11.83$\pm$0.55 16.56$\pm$1.41 29.01$\pm$2.50 31.94$\pm$2.81 28.41$\pm$1.45 33.88$\pm$1.77
LwTR (DA) 19.15$\pm$0.55 23.70$\pm$0.77 11.72$\pm$0.42 16.37$\pm$1.21 15.12$\pm$0.44 20.56$\pm$0.80 11.82$\pm$0.39 15.57$\pm$0.47 28.78$\pm$2.27 31.31$\pm$2.14 28.72$\pm$1.70 34.30$\pm$1.68
MR (DA) 19.13$\pm$0.95 23.17$\pm$1.10 11.59$\pm$0.34 15.89$\pm$0.92 14.66$\pm$1.05 19.63$\pm$1.37 11.50$\pm$0.33 15.62$\pm$0.74 28.65$\pm$3.20 31.23$\pm$3.18 27.08$\pm$1.40 32.39$\pm$1.57
SR (DA) 18.16$\pm$0.63 21.99$\pm$0.91 11.38$\pm$0.44 15.44$\pm$0.96 15.29$\pm$0.96 20.11$\pm$1.14 11.71$\pm$0.25 16.31$\pm$0.57 27.30$\pm$4.37 29.85$\pm$4.54 29.72$\pm$0.91 34.74$\pm$1.05
MELM (DA) 18.59$\pm$0.60 22.67$\pm$0.95 10.75$\pm$0.46 14.11$\pm$0.69 13.94$\pm$0.98 18.50$\pm$1.22 11.28$\pm$0.33 15.43$\pm$0.98 25.71$\pm$1.73 28.19$\pm$1.87 26.58$\pm$1.48 31.47$\pm$1.64
Table 2: Results of existing methods and data augmentation methods in OntoNotes 5.0 in ID setting. The best results are shown in bold. † indicates significantly improved than existing methods (p < 0.05).
5.3 Evaluation Details
The NER model calibration is evaluated based on the "Event of Interests" concept introduced in the previous study Kuleshov and Liang (2015); Jagannatha and Yu (2020). Since the full label space $|\mathcal{Y}|$ is large for structured prediction tasks such as NER, we focus instead on the event set $L(x)$, which is the set containing the events of interest $E\in L(x)$ obtained by processing the model output.
There are two main strategies for constructing $L(x)$: The first strategy is to construct $L(x)$ only from the events obtained by the MAP label sequence prediction of the model; The second strategy is to construct $L(x)$ from all possible label sequences; The first strategy is easy to obtain events, but the coverage of events is low depending on the model’s prediction. The second strategy provides a high coverage of events, but is computationally expensive to obtain events. Jagannatha and Yu (2020) is based on the first strategy, where the entities extracted by the NER model are calibrated on the basis of forecasters (e.g., gradient boosting decision trees Friedman (2000)), which are binary classifiers separate from the NER model. Since the training dataset for forecasters consists of entities extracted by the NER model, more entities are needed to improve the uncertainty performance of the forecasters. Therefore, for example, the top-k Viterbi decoding of the CRF is used to increase the entity coverage and the size of the forecaster’s training dataset.
On the other hand, Jiang et al. (2022) is based on the second strategy, where it introduces a method to find the probability that a span has a specific entity type for datasets with short sequences, such as WikiAnn Pan et al. (2017), with restricted token sequences and span lengths. However, this method is computationally difficult for datasets with longer token sequences and more complex label spaces, such as OntoNotes 5.0 and MultiCoNER, because the number of spans explodes. We therefore simplify the evaluation process by measuring the calibration of the entity span obtained from the NER model’s MAP label sequence prediction of the model. Uncertainty performance is evaluated by taking the product of the probabilities of each token corresponding to an entity as the probability of one entity.
6 Results and Discussion
We present the performance of cross-genre and cross-lingual confidence calibration and uncertainty estimation as the main results. The cross-genre evaluations are quantified by learning on a training set in one genre and evaluating calibration and uncertainty on a test in another genre. Similarly, in the cross-lingual evaluations, we train the model in one language (in this research, we use English; $\mathtt{EN}$) and evaluate the calibration and uncertainty on a test set in another language.
6.1 Cross-genre Evaluation
OntoNotes 5.0 ($\mathtt{bc}$)
Methods $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓)
Baseline 17.54$\pm$0.67 25.90$\pm$1.29 18.83$\pm$0.89 25.65$\pm$1.09 23.52$\pm$0.77 34.25$\pm$1.41 26.20$\pm$1.23 28.76$\pm$1.30 57.47$\pm$0.87 62.96$\pm$0.56
TS 17.19$\pm$0.81 24.93$\pm$1.27 19.42$\pm$1.48 26.32$\pm$1.97 23.51$\pm$1.08 33.68$\pm$1.72 26.85$\pm$2.11 29.36$\pm$2.35 57.66$\pm$1.32 62.96$\pm$1.15
LS 17.45$\pm$0.96 25.43$\pm$1.77 19.38$\pm$1.03 26.36$\pm$1.56 23.72$\pm$1.01 34.23$\pm$1.95 26.34$\pm$1.78 28.81$\pm$2.04 56.98$\pm$1.17 62.51$\pm$0.91
MC Dropout 17.50$\pm$0.66 25.77$\pm$1.58 19.22$\pm$1.21 26.39$\pm$1.16 23.67$\pm$0.73 34.51$\pm$1.59 26.32$\pm$1.10 28.66$\pm$1.12 57.51$\pm$1.29 62.80$\pm$0.90
LwTR (DA) 17.58$\pm$0.44 25.45$\pm$1.34 19.34$\pm$1.34 26.11$\pm$1.56 23.65$\pm$0.53 33.89$\pm$1.13 27.50$\pm$1.73 29.70$\pm$2.01 58.68$\pm$1.51 63.83$\pm$1.22
MR (DA) 17.43$\pm$0.62 24.99$\pm$1.36 18.38$\pm$1.62 24.93$\pm$1.73 23.28$\pm$0.54 33.35$\pm$1.16 26.78$\pm$2.19 28.85$\pm$2.21 59.01$\pm$0.99 64.06$\pm$0.76
SR (DA) 17.01$\pm$0.39 24.45$\pm$0.74 20.01$\pm$1.56 26.94$\pm$1.72 23.42$\pm$0.66 33.29$\pm$1.33 26.62$\pm$1.59 28.81$\pm$1.76 58.14$\pm$0.79 63.02$\pm$0.59
MELM (DA) 17.22$\pm$0.65 24.55$\pm$1.41 19.41$\pm$0.80 26.01$\pm$1.06 23.66$\pm$0.85 33.75$\pm$1.46 30.11$\pm$1.39 32.59$\pm$1.71 58.72$\pm$1.42 63.71$\pm$1.18
OntoNotes 5.0 ($\mathtt{bn}$)
Methods $\mathtt{bc}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓)
Baseline 19.30$\pm$0.82 24.37$\pm$1.47 20.55$\pm$1.59 26.62$\pm$2.55 20.05$\pm$0.98 28.44$\pm$2.25 25.42$\pm$0.73 27.56$\pm$0.64 59.02$\pm$1.16 63.61$\pm$0.66
TS 19.20$\pm$0.88 24.18$\pm$1.75 21.21$\pm$1.14 27.20$\pm$1.72 20.34$\pm$0.73 28.80$\pm$2.12 25.33$\pm$1.28 27.57$\pm$1.27 59.11$\pm$1.06 63.60$\pm$0.60
LS 18.37$\pm$0.60 22.52$\pm$1.41 21.61$\pm$0.47 27.04$\pm$1.04 19.98$\pm$0.41 27.64$\pm$1.11 24.66$\pm$0.48 26.69$\pm$0.44 59.92$\pm$0.75 63.87$\pm$0.77
MC Dropout 18.76$\pm$0.97 23.34$\pm$1.56 20.91$\pm$0.96 26.62$\pm$1.82 20.04$\pm$0.57 28.25$\pm$1.62 25.21$\pm$1.27 27.52$\pm$1.17 59.09$\pm$0.99 63.63$\pm$0.54
LwTR (DA) 20.30$\pm$0.87 25.42$\pm$1.18 20.71$\pm$1.01 27.14$\pm$1.16 20.51$\pm$0.41 29.04$\pm$1.26 26.36$\pm$2.08 28.67$\pm$2.09 59.32$\pm$0.97 64.00$\pm$0.55
MR (DA) 19.78$\pm$1.26 24.35$\pm$1.85 20.19$\pm$0.47 26.08$\pm$1.07 20.42$\pm$0.60 27.83$\pm$1.74 25.69$\pm$0.77 27.75$\pm$0.81 59.57$\pm$0.96 64.13$\pm$0.50
SR (DA) 19.61$\pm$0.97 24.08$\pm$1.64 19.79$\pm$0.75 25.52$\pm$1.22 19.81$\pm$0.39 27.18$\pm$1.30 26.20$\pm$1.56 28.42$\pm$1.68 59.86$\pm$0.67 63.66$\pm$0.40
MELM (DA) 19.93$\pm$0.69 23.98$\pm$1.09 20.40$\pm$0.65 25.54$\pm$1.19 19.73$\pm$0.65 26.80$\pm$1.19 28.47$\pm$2.14 30.59$\pm$2.15 60.51$\pm$0.57 64.44$\pm$0.33
OntoNotes 5.0 ($\mathtt{nw}$)
Methods $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{tc}$ $\mathtt{wb}$
ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓)
Baseline 20.65$\pm$1.79 25.32$\pm$2.15 15.24$\pm$0.65 21.06$\pm$1.21 22.67$\pm$1.24 28.48$\pm$2.17 27.81$\pm$1.26 30.21$\pm$1.39 60.28$\pm$1.17 64.30$\pm$0.86
TS 21.08$\pm$0.75 25.80$\pm$1.01 15.61$\pm$0.46 21.63$\pm$0.80 22.76$\pm$1.01 28.92$\pm$1.47 28.02$\pm$1.61 30.21$\pm$1.90 60.37$\pm$0.89 64.61$\pm$0.68
LS 20.46$\pm$1.23 24.63$\pm$2.21 15.51$\pm$0.55 20.80$\pm$1.70 22.66$\pm$1.10 28.35$\pm$1.86 28.50$\pm$1.52 30.41$\pm$1.21 60.17$\pm$1.05 64.07$\pm$0.72
MC Dropout 21.25$\pm$1.84 25.98$\pm$2.09 15.58$\pm$0.98 21.59$\pm$1.71 22.38$\pm$1.10 28.34$\pm$1.67 28.05$\pm$1.70 30.19$\pm$1.79 60.64$\pm$0.94 64.63$\pm$0.57
LwTR (DA) 21.87$\pm$0.87 26.58$\pm$0.99 15.81$\pm$0.30 21.93$\pm$0.41 22.76$\pm$0.93 28.38$\pm$0.92 27.60$\pm$0.72 29.48$\pm$0.45 59.96$\pm$0.46 64.06$\pm$0.40
MR (DA) 21.70$\pm$0.27 26.29$\pm$0.30 15.55$\pm$0.87 21.38$\pm$2.16 21.08$\pm$1.21 26.33$\pm$2.14 30.35$\pm$2.69 32.44$\pm$2.82 61.16$\pm$1.06 65.12$\pm$0.80
SR (DA) 21.29$\pm$1.37 25.82$\pm$1.31 16.00$\pm$0.58 21.72$\pm$0.22 21.83$\pm$0.67 27.37$\pm$0.85 33.41$\pm$5.50 35.59$\pm$5.44 60.58$\pm$0.72 64.50$\pm$0.54
MELM (DA) 21.96$\pm$1.31 26.91$\pm$1.88 15.83$\pm$0.84 21.76$\pm$1.63 21.16$\pm$1.38 26.88$\pm$1.49 33.92$\pm$4.15 36.39$\pm$4.03 60.94$\pm$0.62 65.03$\pm$0.33
OntoNotes 5.0 ($\mathtt{tc}$)
Methods $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{wb}$
ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓)
Baseline 36.70$\pm$1.65 44.25$\pm$1.66 35.47$\pm$2.48 45.75$\pm$2.46 37.15$\pm$1.77 47.34$\pm$1.79 39.08$\pm$0.56 52.50$\pm$1.41 46.38$\pm$1.28 54.29$\pm$1.37
TS 35.69$\pm$2.21 43.34$\pm$2.18 34.15$\pm$2.65 44.48$\pm$2.56 36.38$\pm$1.79 46.71$\pm$1.43 38.59$\pm$1.53 52.58$\pm$1.38 47.20$\pm$0.92 55.31$\pm$1.10
LS 33.91$\pm$1.86 41.50$\pm$1.75 31.40$\pm$2.35 41.24$\pm$2.43 34.14$\pm$1.91 44.37$\pm$1.42 37.04$\pm$2.25 50.00$\pm$1.92 48.48$\pm$1.29 56.10$\pm$0.89
MC Dropout 35.83$\pm$2.02 43.93$\pm$1.75 33.87$\pm$2.02 44.31$\pm$1.92 36.18$\pm$2.43 46.31$\pm$2.43 38.97$\pm$0.83 52.80$\pm$1.08 46.92$\pm$2.04 54.95$\pm$2.13
LwTR (DA) 34.94$\pm$2.42 43.20$\pm$1.90 32.61$\pm$3.16 43.28$\pm$2.55 34.44$\pm$1.83 44.98$\pm$1.88 37.85$\pm$2.13 52.09$\pm$1.60 46.78$\pm$1.26 54.94$\pm$1.84
MR (DA) 35.18$\pm$2.89 42.62$\pm$2.30 33.50$\pm$3.77 42.66$\pm$3.20 34.35$\pm$2.78 44.78$\pm$2.69 37.97$\pm$2.64 50.85$\pm$3.46 48.61$\pm$1.70 55.78$\pm$1.90
SR (DA) 34.58$\pm$2.40 42.51$\pm$1.55 32.66$\pm$4.13 42.57$\pm$3.28 32.69$\pm$3.21 43.01$\pm$2.83 38.50$\pm$1.51 52.00$\pm$1.56 46.99$\pm$1.27 54.86$\pm$1.40
MELM (DA) 33.05$\pm$1.75 40.55$\pm$2.16 29.46$\pm$1.55 37.81$\pm$1.56 33.46$\pm$1.66 42.78$\pm$2.55 36.79$\pm$1.27 49.33$\pm$2.26 50.52$\pm$1.10 57.27$\pm$1.27
Table 3: Results of existing methods and data augmentation methods in OntoNotes 5.0 in OOD test dataset.
Methods OntoNotes 5.0 ($\mathtt{bc}$) OntoNotes 5.0 ($\mathtt{bn}$)
$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$ $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
Baseline 94.72$\pm$0.21 95.13$\pm$0.43 96.40$\pm$0.40 93.27$\pm$0.41 92.69$\pm$0.57 93.03$\pm$0.56 95.12$\pm$0.30 97.23$\pm$0.20 95.83$\pm$0.45 95.29$\pm$0.27 93.62$\pm$0.59 93.13$\pm$0.40
TS 94.89$\pm$0.59 95.14$\pm$0.35 96.15$\pm$0.51 93.26$\pm$0.45 92.78$\pm$1.01 92.97$\pm$0.83 95.05$\pm$0.39 97.38$\pm$0.17 95.33$\pm$0.31 95.23$\pm$0.20 93.96$\pm$0.51 93.25$\pm$0.29
LS 94.74$\pm$0.54 95.09$\pm$0.37 96.15$\pm$0.30 93.15$\pm$0.43 92.60$\pm$0.79 92.73$\pm$0.36 94.99$\pm$0.22 97.32$\pm$0.20 95.60$\pm$0.22 95.11$\pm$0.37 93.49$\pm$0.43 92.90$\pm$0.47
MC Dropout 94.71$\pm$0.31 95.09$\pm$0.18 96.07$\pm$0.24 93.11$\pm$0.43 92.76$\pm$0.67 92.88$\pm$0.33 95.03$\pm$0.34 97.30$\pm$0.18 95.78$\pm$0.46 95.29$\pm$0.19 93.80$\pm$0.44 93.22$\pm$0.35
LwTR (DA) 94.53$\pm$0.28 95.02$\pm$0.37 96.22$\pm$0.33 93.23$\pm$0.23 92.76$\pm$0.64 92.91$\pm$0.52 94.36$\pm$0.54 97.29$\pm$0.14 95.74$\pm$0.16 95.15$\pm$0.20 93.64$\pm$0.51 93.08$\pm$0.49
MR (DA) 94.44$\pm$0.29 94.88$\pm$0.24 96.53$\pm$0.43 93.4$\pm$0.29 92.82$\pm$0.60 92.74$\pm$0.42 94.57$\pm$0.50 97.20$\pm$0.19 96.27$\pm$0.31 95.11$\pm$0.22 93.64$\pm$0.55 92.91$\pm$0.52
SR (DA) 94.44$\pm$0.35 95.09$\pm$0.32 95.70$\pm$0.40 93.21$\pm$0.37 93.24$\pm$0.43 93.06$\pm$0.39 94.76$\pm$0.65 97.28$\pm$0.15 95.85$\pm$0.33 95.30$\pm$0.17 93.78$\pm$0.63 93.06$\pm$0.24
MELM (DA) 94.51$\pm$0.16 95.15$\pm$0.34 96.01$\pm$0.29 93.09$\pm$0.44 92.64$\pm$0.52 92.90$\pm$0.47 94.34$\pm$0.47 97.24$\pm$0.21 96.18$\pm$0.32 95.32$\pm$0.32 93.51$\pm$0.50 92.97$\pm$0.48
Methods OntoNotes 5.0 ($\mathtt{nw}$) OntoNotes 5.0 ($\mathtt{tc}$)
$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$ $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
Baseline 94.60$\pm$0.80 96.36$\pm$0.32 95.22$\pm$0.48 97.81$\pm$0.12 93.32$\pm$0.44 93.29$\pm$0.46 87.10$\pm$1.25 89.22$\pm$0.71 84.94$\pm$1.61 81.28$\pm$2.58 93.45$\pm$0.77 89.62$\pm$1.10
TS 94.50$\pm$0.40 96.36$\pm$0.32 95.34$\pm$0.39 97.74$\pm$0.18 93.15$\pm$0.52 93.33$\pm$0.37 87.74$\pm$1.12 89.45$\pm$0.47 85.95$\pm$1.65 82.50$\pm$1.35 93.11$\pm$0.98 89.93$\pm$0.88
LS 94.65$\pm$0.30 96.23$\pm$0.24 95.19$\pm$0.57 97.70$\pm$0.09 93.05$\pm$0.43 93.39$\pm$0.41 87.07$\pm$1.00 89.57$\pm$0.76 86.67$\pm$1.75 82.79$\pm$1.09 92.75$\pm$1.06 90.66$\pm$0.61
MC Dropout 94.37$\pm$0.92 96.32$\pm$0.23 95.27$\pm$0.31 97.81$\pm$0.24 93.40$\pm$0.25 93.15$\pm$0.47 87.25$\pm$0.73 89.02$\pm$1.08 85.12$\pm$1.62 81.95$\pm$2.56 93.36$\pm$0.89 90.05$\pm$0.84
LwTR (DA) 94.11$\pm$0.68 96.33$\pm$0.22 95.36$\pm$0.29 97.79$\pm$0.31 94.11$\pm$0.27 92.76$\pm$0.25 86.95$\pm$0.61 89.74$\pm$0.72 86.20$\pm$1.67 83.08$\pm$1.78 93.70$\pm$0.64 90.28$\pm$0.55
MR (DA) 93.43$\pm$0.13 96.18$\pm$0.33 95.01$\pm$0.69 97.69$\pm$0.12 93.15$\pm$0.60 92.67$\pm$0.32 86.78$\pm$1.12 90.06$\pm$0.61 86.36$\pm$1.64 83.81$\pm$2.79 93.69$\pm$0.61 90.69$\pm$1.23
SR (DA) 94.18$\pm$0.92 96.21$\pm$0.30 95.45$\pm$0.30 97.87$\pm$0.14 93.41$\pm$0.23 93.39$\pm$0.29 86.78$\pm$1.49 89.61$\pm$0.56 86.42$\pm$2.36 81.83$\pm$2.85 93.53$\pm$0.72 90.04$\pm$0.97
MELM (DA) 94.07$\pm$0.67 96.09$\pm$0.14 95.67$\pm$0.71 97.83$\pm$0.12 92.84$\pm$0.73 93.43$\pm$0.64 86.38$\pm$1.16 89.05$\pm$1.18 86.65$\pm$1.37 81.89$\pm$2.77 93.30$\pm$0.59 89.12$\pm$1.47
Table 4: AUPRC scores of existing methods and data augmentation methods in OntoNotes 5.0.
[…]
Figure 1: t-SNE plot of token embeddings of OntoNotes 5.0 $\mathtt{bn}$ training set (red), generated data by MELM (blue), source domain test set (green) and OntoNotes 5.0 $\mathtt{wb}$ test set (purple), respectively.
We presume that the augmented data is not far from the original training set, because data augmentation methods we targeted in this study are based on the replacement of tokens or entities. Considering a recent study that indicates models tend to be more overconfident in areas with less training data Xiong et al. (2023), we can consider calibration performance in OOD sets, especially far from the source domain, will not improve by data augmentation for NER, while the performance in ID sets will be better than existing methods.
To illustrate this, we performed t-SNE van der Maaten and Hinton (2008) for the token embeddings with only entity token from trained Baseline model, shown in Figure 1. We can understand that the token embeddings from augmented data are near the train set or ID test set, while the OOD test sets have some poorly covered regions. Generating sentences that are distant from the training data set and semantically aligned entities from label description for uncertainty estimation is an interesting direction for future research.
AUPRC scores are shown in Table 4. Among the existing methods, TS shows superior performance; in data augmentation methods, MELM is not as good as in the case of calibration metrics such as ECE and MCE, and MR tends to show superior uncertainty performance.
6.2 Cross-lingual Evaluation
Methods MultiCoNER ($\mathtt{EN}$)
$\mathtt{EN}$ $\mathtt{DE}$ $\mathtt{ES}$ $\mathtt{HI}$
ECE (↓) MCE (↓) AUPRC (↑) ECE (↓) MCE (↓) AUPRC (↑) ECE (↓) MCE (↓) AUPRC (↑) ECE (↓) MCE (↓) AUPRC (↑)
Baseline 28.29$\pm$0.30 30.51$\pm$0.39 93.04$\pm$0.18 31.31$\pm$0.52 34.91$\pm$0.83 91.97$\pm$0.23 31.22$\pm$0.28 33.70$\pm$0.39 90.87$\pm$0.27 46.84$\pm$1.64 48.13$\pm$1.51 82.04$\pm$2.24
TS 28.46$\pm$0.43 30.70$\pm$0.52 93.13$\pm$0.17 31.45$\pm$0.70 35.08$\pm$1.05 92.02$\pm$0.24 31.24$\pm$0.41 33.77$\pm$0.38 90.92$\pm$0.18 46.83$\pm$1.38 48.35$\pm$1.25 83.01$\pm$1.45
LS 28.50$\pm$0.57 30.60$\pm$0.68 93.12$\pm$0.13 31.50$\pm$0.64 34.81$\pm$0.66 91.93$\pm$0.26 31.43$\pm$0.58 33.83$\pm$0.67 90.82$\pm$0.10 46.36$\pm$1.23 47.95$\pm$1.03 84.00$\pm$1.60
MC Dropout 28.57$\pm$0.34 30.83$\pm$0.54 92.97$\pm$0.34 31.64$\pm$0.48 35.24$\pm$0.68 91.86$\pm$0.37 31.47$\pm$0.42 33.98$\pm$0.40 90.79$\pm$0.22 47.42$\pm$1.30 48.77$\pm$1.23 81.39$\pm$3.30
LwTR (DA) 28.17$\pm$0.54 30.48$\pm$0.77 92.80$\pm$0.28 31.13$\pm$0.59 34.60$\pm$0.78 91.57$\pm$0.34 31.10$\pm$0.35 33.61$\pm$0.51 90.66$\pm$0.27 46.70$\pm$1.47 47.95$\pm$1.30 82.57$\pm$1.96
MR (DA) 28.01$\pm$0.42 30.08$\pm$0.49 93.30$\pm$0.24 31.12$\pm$0.74 34.71$\pm$0.81 92.05$\pm$0.20 30.75$\pm$0.34 33.24$\pm$0.36 91.03$\pm$0.15 46.96$\pm$1.20 48.28$\pm$1.12 81.75$\pm$2.52
SR (DA) 28.15$\pm$0.42 30.36$\pm$0.48 93.08$\pm$0.26 31.17$\pm$0.39 34.42$\pm$0.70 92.02$\pm$0.39 31.60$\pm$0.55 33.86$\pm$0.56 90.65$\pm$0.33 45.85$\pm$0.53 47.38$\pm$0.47 84.91$\pm$0.91
MELM (DA) 28.53$\pm$0.38 30.68$\pm$0.43 92.72$\pm$0.22 32.61$\pm$0.49 36.14$\pm$0.65 91.17$\pm$0.29 32.09$\pm$0.44 34.38$\pm$0.52 90.14$\pm$0.30 47.91$\pm$1.79 49.18$\pm$1.79 81.13$\pm$2.41
Table 5: Results of existing calibration methods and data augmentation methods in MultiCoNER.
The results of cross-lingual transfer in MultiCoNER are shown in Table 5 with English as the source language. MR performs better in uncertainty performance for the ID situation. In contrast to the calibration and uncertainty performance in the cross-genre setting, both MR and SR show better calibration and uncertainty in the OOD setting. In Jiang et al. (2022), the result shows that the larger the linguistic distance Chiswick and Miller (2005), the more lenient the calibration and uncertainty estimation tends to be, and similar trends are obtained in this experiment. Unlike the discussion in Section 6.1, the uncertainty performance by data augmentation is also good for OOD in cross-lingual setting because the areas where only target set exist is limited in MultiCoNER (illustrated in Appendix F). On the other hand, MELM, which tends to show excellent calibration performance in cross-genre calibration, does not show good performance in cross-lingual settings.
The amount of data for each language in the CC100 Conneau et al. (2020) dataset used to train the base model, mDeBERTaV3, was highest for English, followed by German, Spanish, Hindi, and Bangla which correlates with the trend of the calibration results. Moreover, as mentioned in Limisiewicz et al. (2023), languages that tend to have vocabulary overlap between languages in tokenization perform better in cross-lingual transfer in NER. Similar effects may be observed in confidence calibration and uncertainty estimation.
6.3 Detailed Analyzes
We investigate the effects of entity overlap rates and the perplexity of the generated sentences to gain a better understanding of the confidence calibration and uncertainty estimation performance of data augmentation methods for NER. We also investigate the impact of data augmentation size in several settings.
Figure 2: Average values of evaluation metrics for each data augmentation method in ID settings.
Figure 3: Average values of evaluation metrics for each data augmentation method in OOD settings.
Methods OntoNotes 5.0 ($\mathtt{bc}$) OntoNotes 5.0 ($\mathtt{bn}$) OntoNotes 5.0 ($\mathtt{nw}$) OntoNotes 5.0 ($\mathtt{tc}$) MultiCoNER ($\mathtt{EN}$)
LwTR 7.05 7.59 8.28 7.33 6.78
MR 5.36 5.27 5.27 5.83 5.83
SR 5.91 6.35 6.62 6.02 6.35
MELM 5.56 5.65 5.55 5.90 6.14
(Train) 5.18 4.84 4.86 5.80 5.54
Table 6: Sentences perplexities generated by the data augmentation method in each dataset. Each data augmentation method is performed to increase the training data. Bold means the lowest score in data augmentation methods.
6.3.1 Impact of Augmentation Size
To investigate the impact of data augmentation size on calibration and uncertainty performance, we analyze the trend of evaluation metrics in $\mathtt{tc}$$\mathtt{mz}$ scenario of OntoNotes 5.0 and $\mathtt{EN}$$\mathtt{ES}$ scenario of MultiCoNER, respectively. Figure 2 and 3 illustrate the results in the ID and OOD settings, respectively. In many cases, MR improves the calibration and uncertainty performance by increasing data. SR consistently improves as the dataset size doubles, whereas LwTR demonstrates only marginal improvement or even worsens as the dataset size increases. Finally, MELM improves further for OntoNotes 5.0 $\mathtt{tc}$, which shows excellent performance, and deteriorates further for MultiCoNER $\mathtt{EN}$, which shows poor performance.
These results show that the calibration algorithm with the best performance for cross-domain transfers is likely to have better performance as the augmentation size is increased. On the other hand, increasing the augmentation size in MR improves the calibration and uncertainty performance compared to similar other data augmentation methods.
Since data augmentation by MR and MELM is performed only on the entity region, the uncertainty estimation performance is relatively less adversely affected by increasing the data augmentation size. On the other hand, in SR and LwTR, data augmentation that replaces tokens may often inject tokens with inappropriate parts of speech for that sentence, so increasing the data augmentation size often leads to a degradation of uncertainty estimation performance.
6.3.2 Impact of Perplexities for Augmented Sentences
To investigate the influence of replacement units on data augmentation for NER as mentioned in Section 6.3.1, we measured the perplexity of the augmented sentences using GPT-2 Radford et al. (2019). The average perplexities of the augmented sentences and the average perplexities of the original training set for each dataset are shown in Table 6. Lower perplexity from augmented sentences tends to improve calibration performance and uncertainty performance. Consistently, the average perplexity of the sentences generated by MR is the lowest. Since MR performs substitutions on an entity-by-entity basis and does not affect the structure of the sentence itself, it has the lowest perplexity among the data augmentation methods in NER. MELM has the second lowest perplexity after MR, and may be adversely affected by generated entities that are adapted to the context but not actually present.
7 Conclusion
In this paper, we investigated the impact of data augmentation on the confidence calibration and uncertainty estimation in NER in terms of genre and language, using several metrics. First, we find that MELM, MR, and SR lead to better calibration and uncertainty performance in the ID setting consistently. On the other hand, in the OOD setting, uncertainty estimation by data augmentation is less effective, especially when the target domain is far from the source domain. Second, our results suggest that the lower the perplexity of the augmented data, as in MR, the further better the calibration and uncertainty performance as the augmentation size is increased. Data augmentation methods for NER do not require changes to the model structure and only require more data to improve entity-level calibration and performance without the need to change the model structure. Our findings indicate the effectiveness of uncertainty estimation through data augmentation for NER, and will be expected to stimulate future research based on their limitations.
Limitations
While this experiment provided valuable insights into the impact of data augmentation on confidence calibration and uncertainty estimation in NER across different genres and languages, there are several limitations that should be acknowledged.
First, due to resource limitations, the experiment was limited to evaluation with English as the source language. To effectively investigate the calibration and uncertainty of zero-shot cross-lingual transfer, it is important to expand the investigation to include a wider range of languages as the source language. Therefore, future research should prioritize the investigation of calibration and uncertainty performance using different languages as the source for zero-shot cross-lingual transfer.
Second, as mentioned in Section 5.3, regarding the calibration and uncertainty evaluation policy, we simply evaluated an entity span as a single data instance, but a rigorous evaluation method that performs evaluation while considering multiple span candidates has been proposed Jiang et al. (2022). Establishing span-level NER calibration evaluation methods that can efficiently and comprehensively evaluate calibration and uncertainty for entity types for datasets with many entity types and long sequence lengths is a topic for future research.
Lastly, we broadly evaluated the calibration and uncertainty performance in both cross-genre and cross-lingual settings on data augmentation for NER, but only using sequence labeling-based methods. Recently, other paradigms in NER, such as span-based methods Fu et al. (2021) and Seq2Seq (sequence-to-sequence)-based methods Yan et al. (2021), have been proposed. In the future, the calibration or uncertainty performance of these methods could be evaluated.
Ethical Considerations
In this study, we used existing datasets that have cleared ethical issues. Furthermore, the data augmentation methods we used for uncertainty estimation are substitution-based methods except for MELM, and MELM generated entities from existing datasets that have no ethical issues. Therefore, it is unlikely that toxic sentences would be generated.
Acknowledgements
The authors also acknowledge the Nara Institute of Science and Technology’s HPC resources made available for conducting the research reported in this paper.
Appendix A Inference Time
Methods Inference time [s]
Baseline 14.90 $\pm$ 0.10
TS 15.53 $\pm$ 0.92
LS 14.94 $\pm$ 0.24
MC Dropout 271.77 $\pm$ 1.81
LwTR 14.91 $\pm$ 0.10
MR 14.93 $\pm$ 0.17
SR 14.83 $\pm$ 0.12
MELM 14.89 $\pm$ 0.14
Table 7: Inference time for each algorithm on MultiCoNER $\mathtt{EN}$ full test data.
Methods OntoNotes 5.0 ($\mathtt{bc}$) OntoNotes 5.0 ($\mathtt{bn}$)
$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$ $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
Baseline 81.39$\pm$0.78 80.86$\pm$1.03 81.61$\pm$1.36 75.49$\pm$0.90 68.83$\pm$1.27 45.74$\pm$0.74 80.74$\pm$1.21 90.25$\pm$0.36 81.47$\pm$0.96 81.04$\pm$0.64 72.36$\pm$1.88 46.86$\pm$0.52
TS 81.10$\pm$0.94 81.19$\pm$0.89 80.80$\pm$1.37 75.14$\pm$1.60 69.20$\pm$2.73 45.58$\pm$1.02 81.31$\pm$1.18 90.37$\pm$0.49 80.96$\pm$1.32 81.13$\pm$0.62 71.83$\pm$1.76 46.50$\pm$0.69
LS 81.21$\pm$1.11 81.17$\pm$0.91 81.43$\pm$1.33 75.30$\pm$1.26 69.64$\pm$1.45 45.75$\pm$0.82 82.08$\pm$0.62 90.32$\pm$0.36 81.22$\pm$0.52 80.95$\pm$0.37 72.45$\pm$1.38 46.69$\pm$0.60
MC Dropout 81.49$\pm$0.80 81.06$\pm$0.71 81.12$\pm$0.63 75.24$\pm$1.02 69.53$\pm$1.78 45.73$\pm$0.46 81.55$\pm$0.63 90.21$\pm$0.36 80.80$\pm$1.10 81.11$\pm$0.46 73.13$\pm$1.97 46.71$\pm$0.60
LwTR (DA) 80.85$\pm$0.82 80.91$\pm$0.93 81.45$\pm$1.08 75.33$\pm$0.82 68.40$\pm$0.94 45.53$\pm$0.84 79.43$\pm$1.13 89.98$\pm$0.40 80.75$\pm$0.67 80.33$\pm$0.31 69.62$\pm$1.80 46.23$\pm$0.54
MR (DA) 80.93$\pm$0.61 80.88$\pm$0.61 82.02$\pm$0.66 75.66$\pm$0.79 69.49$\pm$1.78 45.38$\pm$0.72 79.93$\pm$1.43 90.07$\pm$0.23 81.70$\pm$0.61 80.54$\pm$0.50 72.44$\pm$1.46 46.45$\pm$0.47
SR (DA) 81.52$\pm$0.69 81.20$\pm$0.78 79.93$\pm$0.95 75.08$\pm$0.89 69.86$\pm$1.30 46.04$\pm$0.57 80.24$\pm$1.44 90.05$\pm$0.21 80.92$\pm$0.93 80.84$\pm$0.42 70.80$\pm$1.66 46.98$\pm$0.61
MELM (DA) 81.08$\pm$0.37 80.81$\pm$0.97 80.11$\pm$0.98 74.74$\pm$1.24 66.68$\pm$1.18 45.19$\pm$1.05 79.23$\pm$0.64 90.26$\pm$0.38 81.48$\pm$0.65 80.66$\pm$0.79 68.42$\pm$1.65 46.36$\pm$0.44
Methods OntoNotes 5.0 ($\mathtt{nw}$) OntoNotes 5.0 ($\mathtt{tc}$)
$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$ $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
Baseline 74.34$\pm$4.10 83.08$\pm$1.19 73.56$\pm$3.31 90.08$\pm$0.31 72.59$\pm$1.34 46.47$\pm$0.59 55.29$\pm$2.01 59.13$\pm$2.80 50.68$\pm$3.51 46.14$\pm$4.31 69.52$\pm$1.45 40.85$\pm$1.36
TS 75.34$\pm$1.67 83.02$\pm$0.98 75.01$\pm$2.21 90.04$\pm$0.24 71.98$\pm$1.17 46.29$\pm$0.87 56.81$\pm$2.05 59.04$\pm$2.95 52.98$\pm$3.34 48.85$\pm$3.26 67.45$\pm$2.30 41.12$\pm$1.27
LS 76.60$\pm$1.65 83.27$\pm$1.49 75.79$\pm$2.00 90.20$\pm$0.26 71.91$\pm$2.67 46.68$\pm$0.69 53.98$\pm$3.40 56.12$\pm$6.02 51.17$\pm$5.94 48.62$\pm$4.82 66.01$\pm$3.26 40.63$\pm$1.83
MC Dropout 75.07$\pm$2.84 82.69$\pm$2.11 73.79$\pm$2.23 89.98$\pm$0.56 71.96$\pm$1.43 46.25$\pm$0.92 55.16$\pm$1.70 58.95$\pm$2.87 51.11$\pm$3.75 47.31$\pm$4.48 69.15$\pm$3.05 40.57$\pm$1.44
LwTR (DA) 74.80$\pm$1.57 83.01$\pm$0.41 75.01$\pm$3.35 89.79$\pm$0.28 70.85$\pm$1.13 46.78$\pm$0.54 54.01$\pm$2.14 60.86$\pm$2.89 53.89$\pm$3.76 50.20$\pm$3.77 69.53$\pm$1.60 40.80$\pm$0.97
MR (DA) 73.57$\pm$1.09 81.52$\pm$2.09 71.43$\pm$3.80 89.90$\pm$0.34 68.31$\pm$3.52 44.88$\pm$1.38 53.73$\pm$2.35 57.46$\pm$3.70 52.74$\pm$3.27 46.90$\pm$4.87 68.57$\pm$2.71 40.50$\pm$1.79
SR (DA) 73.64$\pm$3.45 82.03$\pm$2.14 72.25$\pm$4.88 90.24$\pm$0.11 66.18$\pm$4.59 46.38$\pm$1.45 53.41$\pm$2.46 58.54$\pm$3.20 53.08$\pm$4.85 46.48$\pm$7.08 68.13$\pm$1.41 41.20$\pm$1.23
MELM (DA) 73.46$\pm$2.46 82.22$\pm$1.23 75.56$\pm$2.60 89.94$\pm$0.18 62.43$\pm$2.95 45.19$\pm$0.97 48.01$\pm$5.27 49.59$\pm$6.16 48.93$\pm$4.11 42.09$\pm$5.61 63.46$\pm$2.28 36.16$\pm$3.76
Table 8: F1 scores of existing calibration methods and data augmentation methods in OntoNotes 5.0.
Appendix B Details of Existing Calibration Methods
In this section, we describe the popular baseline methods for confidence calibration. We use the following notations: $\bm{z}_{i}$ denotes the logits for class $i$, $p_{i}$ denotes the calibrated probability for class $i$, $y_{i}$ denotes the label for class $i$, and $K$ denotes the number of classes.
B.1 Temperature Scaling (TS)
TS Guo et al. (2017) is a post-processing technique for calibrating the confidence scores outputted by a neural network. It involves scaling the logits (i.e., the outputs of the final layer before the softmax) by a temperature parameter $T$ before applying the softmax function to obtain the calibrated probabilities. The softmax function takes a vector of logits $\bm{z}$ and returns a distribution $\bm{p}$:
$p_{i}=\frac{\exp(\bm{z}_{i}/T)}{\sum_{j=1}^{K}\exp(\bm{z}_{j}/T)}.$
B.2 Label Smoothing (LS)
LS Miller et al. (1996); Pereyra et al. (2017) is a regularization technique used to improve the calibration and generalization performance of the model. By introducing a small degree of uncertainty in the target labels during training, label smoothing mitigates overfitting and encourages the model to learn more robust and accurate representations, ultimately contributing to improved overall performance on the task at hand. LS is characterized by introducing a smoothing parameter $\epsilon$ and smoothed label $y^{LS}_{i}$ as follows,
$y^{LS}_{i}=y_{i}(1-\epsilon)+\frac{\epsilon}{K}.$
B.3 Monte-Carlo Dropout (MC Dropout)
MC Dropout is a regularization technique that can be used for uncertainty estimation in neural networks Gal and Ghahramani (2016). In this method, we need to run the model $M$ times with different dropout masks and take the average softmax output over all the runs (We use $M=20$). The procedure can be represented using the following formula:
$p_{i}=\frac{1}{M}\sum_{t=1}^{M}\frac{\exp(\bm{z}_{i}^{(t)})}{\sum_{j=1}^{K}\exp(\bm{z}_{j}^{(t)})}.$
Methods $\mathtt{EN}$ $\mathtt{DE}$ $\mathtt{ES}$ $\mathtt{HI}$
Baseline 68.80$\pm$0.38 64.91$\pm$0.60 63.53$\pm$0.41 37.33$\pm$3.77
TS 68.51$\pm$0.52 64.70$\pm$0.90 63.41$\pm$0.45 37.90$\pm$2.79
LS 69.17$\pm$0.55 65.37$\pm$0.51 63.83$\pm$0.32 39.93$\pm$3.50
MC Dropout 68.56$\pm$0.96 64.70$\pm$0.87 63.39$\pm$0.69 36.38$\pm$5.89
LwTR (DA) 68.86$\pm$0.82 64.95$\pm$0.64 63.52$\pm$0.85 38.24$\pm$3.11
MR (DA) 69.71$\pm$0.72 65.37$\pm$0.57 64.25$\pm$0.62 37.53$\pm$4.03
SR (DA) 68.81$\pm$0.41 64.75$\pm$0.86 63.85$\pm$0.46 42.31$\pm$1.45
MELM (DA) 68.57$\pm$0.54 63.40$\pm$0.49 62.76$\pm$0.64 37.78$\pm$3.16
Table 9: F1 scores of existing calibration methods and data augmentation methods in MultiCoNER.
Appendix C F1 Scores
Table 8 and 9 show F1 scores. Note that in many cases, data augmentation methods do not degrade predictive performance itself, but MELM often significantly degrades predictive performance in some cases, especially when the source domains are $\mathtt{nw}$ and $\mathtt{tc}$. Considering Section 6.1 and 6.2, MR improves calibration and uncertainty performance in many cases without degrading predictive performance.
Appendix D More Results about Test Set Duplication
Table 10 shows the results of the percentage increase in entity duplication that are new overlaps with each target domain’s test set when applying each data augmentation method except MR, where the source domains are $\mathtt{bc}$, $\mathtt{bn}$, and $\mathtt{nw}$. In all cases there is only a small increase. These results and the MR, which shows good calibration and uncertainty performance indicated from Section 6.1 and 6.2, do not increase the number of new entities in the training data set suggest that the entity overlap rate does not affect calibration and uncertainty estimation.
Methods OntoNotes 5.0 ($\mathtt{bc}$)
$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
LwTR 0.27 0.26 0.00 0.14 1.83 0.30
SR 0.00 0.18 0.00 0.14 0.00 0.15
MELM 0.41 0.53 0.19 0.17 0.91 0.45
Methods OntoNotes 5.0 ($\mathtt{bn}$)
$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
LwTR 0.55 0.35 0.19 0.35 0.91 0.60
SR 0.55 0.26 0.19 0.21 0.00 0.45
MELM 0.68 0.35 0.37 0.10 0.46 0.30
Methods OntoNotes 5.0 ($\mathtt{nw}$)
$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
LwTR 0.96 1.23 0.37 0.52 5.02 1.34
SR 0.41 0.09 0.56 0.21 0.46 1.04
MELM 1.10 0.79 1.48 0.55 1.37 0.45
Table 10: The percentage of new entities generated by each data augmentation method using the training set in the case of the source domain $\mathtt{bc}$, $\mathtt{bn}$ and $\mathtt{nw}$.
Methods OntoNotes 5.0 ($\mathtt{bc}$) OntoNotes 5.0 ($\mathtt{bn}$) OntoNotes 5.0 ($\mathtt{nw}$) OntoNotes 5.0 ($\mathtt{tc}$) MultiCoNER ($\mathtt{EN}$)
LwTR 27.77 32.69 38.65 19.83 18.46
MR 0.00 0.00 0.00 0.00 0.00
SR 25.23 26.34 35.13 8.56 20.45
MELM 45.26 45.95 43.37 34.75 37.64
Table 11: The percentage increase in new entities when each data augmentation method is performed on the original train set.
Methods $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
LwTR 0.00 0.00 0.00 0.10 0.00 0.00
SR 0.14 0.00 0.00 0.10 0.00 0.15
MELM 0.27 0.35 0.19 0.14 0.00 0.30
Table 12: The percentage increase in entity duplication in the case of the source domain $\mathtt{tc}$ that are new overlaps with each target domain’s test set when applying each data augmentation method except MR. More results are in Appendix D.
Appendix E Impact of New Entities via Data Augmentation
To investigate the impact of new entities added by data augmentation methods on calibration performance, we measured the percentage of new entities added in the training data and the percentage of new entities that overlap with the test set. Table 11 shows the percentage of new entities increased by data augmentation with the train set as the source domain in each dataset. In all data sets, MELM has observed the most increase of the new entities in the augmented data set. On the other hand, MR that shows good calibration performance followed by MELM does not increase the number of new entities because the replacement is based on the entities in the original training data. Furthermore, the entities generated have little overlap with the target domain, as shown in Table 12. Therefore, new entities by data augmentation methods for NER are likely to have no effect on calibration performance or uncertainty performance.
Appendix F t-SNE Plot for MultiCoNER Dataset
To overview of the ID and OOD data instances in the MultiCoNER dataset, t-SNE plot is shown in Figure 4.
Appendix G Results for Low-resource Language
To investigate the uncertainty estimation performance for low-resource language, we additionaly show the results of 10,000 examples of Bangla ($\mathtt{BN}$) from MultiCoNER dataset in Table 13 when source language is $\mathtt{EN}$. The results show that data augmentation is also effective in uncertainty estimation for low-resource language.
Figure 4: t-SNE plot of token embeddings of MultiCoNER $\mathtt{EN}$ training set (red), generated data by SR (blue), source domain test set (green) and MultiCoNER $\mathtt{HI}$ test set (purple), respectively.
Methods ECE (↓) MCE (↓) AUPRC (↑)
Baseline 49.60$\pm$2.02 51.32$\pm$1.96 79.49$\pm$2.21
TS 48.85$\pm$1.89 50.60$\pm$1.60 79.09$\pm$4.22
LS 48.00$\pm$1.97 49.91$\pm$1.54 79.60$\pm$3.51
MC Dropout 49.29$\pm$2.20 50.93$\pm$2.14 78.31$\pm$2.52
LwTR (DA) 48.66$\pm$1.35 50.22$\pm$1.36 80.93$\pm$1.75
MR (DA) 49.54$\pm$2.65 51.20$\pm$2.65 79.17$\pm$2.97
SR (DA) 47.67$\pm$0.98 49.46$\pm$0.88 81.96$\pm$1.35
MELM (DA) 50.77$\pm$0.88 52.15$\pm$0.81 75.55$\pm$2.59
Table 13: Results of existing methods and data augmentation methods in MultiCoNER $\mathtt{BN}$.
Appendix H Licenses of Datasets
OntoNotes 5.0 can be used for research purposes as described in https://catalog.ldc.upenn.edu/LDC2013T19. MultiCoNER dataset is licensed by CC BY 4.0 as described in https://aws.amazon.com/marketplace/pp/prodview-cdhrtt7vq4hf4.

### v2

@@
 The improvement tends to be better the lower the perplexity of the sentences generated by the data augmentation.
-Our code will be released after acceptance.
+Our code is available on https://github.com/wataruhashimoto52/ner_da_uncertainty.
 2 Related Work
@@
 For MultiCoNER, we randomly sampled and fixed 10,000 cases out of 200,000 test cases.
+Dataset &\&& Domain LwTR MR SR MELM ($\eta$, $\mu$)
+OntoNotes 5.0
+$\mathtt{bc}$ 0.3 0.7 0.3 (0.5, 0.5)
+$\mathtt{bn}$ 0.4 0.8 0.2 (0.7, 0.3)
+$\mathtt{mz}$ 0.7 0.4 0.5 (0.3, 0.3)
+$\mathtt{nw}$ 0.7 0.5 0.7 (0.7, 0.7)
+$\mathtt{tc}$ 0.4 0.4 0.1 (0.3, 0.3)
+$\mathtt{wb}$ 0.7 0.7 0.8 (0.5, 0.7)
+MultiCoNER
+English ($\mathtt{EN}$) 0.2 0.8 0.4 (0.3, 0.3)
+Table 2: Optimized hyperparameters in data augmentation methods in each source domain. We present the binomial distribution parameters for LwTR, SR and MR, and ($\eta$, $\mu$) for MELM, respectively.
 5.2 Training Details
@@
 (2019) to optimize the temperature parameter based on dev set loss with a search range of [0.001, 0.002, …, 5.000] in 100 trials.
-In addition, we optimized the binomial distribution parameter for data augmentation methods using the dev set by a grid search in the range of [0.1, 0.2, …, 0.8].
+In addition, we optimized the binomial distribution parameter to manipulate replacement intensity for data augmentation methods using the dev set by a grid search in the range of [0.1, 0.2, …, 0.8].
 In LS, we conducted a grid search in the range of [0.01, 0.05, 0.1, 0.2, 0.3] to optimize the smoothing parameter.
@@
 (2022).
+All hyperparameters in data augmentation are shown in Table  2.
+The implementations of LwTR, MR and SR are used several repos, [^2] https://github.com/boschresearch/data-augmentation-coling2020  [^3] https://github.com/kajyuuen/daaja while the implementation of MELM used the official repo.
+[^4] https://github.com/RandyZhouRan/MELM
 We perform each experiment 10 times using different random seeds, collect evaluation metric values, and report their average and standard deviation.
@@
 The best results are shown in bold.
-† indicates significantly improved than existing methods (p < 0.05).
+† indicates significantly improved than existing methods (p < 0.05) by using t-test.
 5.3 Evaluation Details
@@
 AUPRC scores are shown in Table 5.
+In the AUPRC scores in OntoNotes 5.0, data augmentation methods are outperform existing methods in 15 cases out of 24 cases.
 Among the existing methods, TS shows superior performance; in data augmentation methods, MELM is not as good as in the case of calibration metrics such as ECE and MCE, and MR tends to show superior uncertainty performance.
+Calibration and scores based on AUC measure different points of uncertainty Galil et al.
+(2023), therefore we assume that uncertainties that can be improved vary depending on the methods.
 6.2 Cross-lingual Evaluation
@@
 (2022), the result shows that the larger the linguistic distance Chiswick and Miller (2005), the more lenient the calibration and uncertainty estimation tends to be, and similar trends are obtained in this experiment.
-Unlike the discussion in Section 6.1, the uncertainty performance by data augmentation is also good for OOD in cross-lingual setting because the areas where only target set exist is limited in MultiCoNER (illustrated in Appendix F).
+Unlike the discussion in Section 6.1, the uncertainty performance by data augmentation is also good for OOD in cross-lingual setting because the areas where only target set exist is limited in MultiCoNER (illustrated in Appendix G).
 On the other hand, MELM, which tends to show excellent calibration performance in cross-genre calibration, does not show good performance in cross-lingual settings.
@@
 In many cases, MR improves the calibration and uncertainty performance by increasing data.
+[^5] Note that we have not discussed about the absolute values of the uncertainty estimation performance.
 SR consistently improves as the dataset size doubles, whereas LwTR demonstrates only marginal improvement or even worsens as the dataset size increases.
@@
 Since MR performs substitutions on an entity-by-entity basis and does not affect the structure of the sentence itself, it has the lowest perplexity among the data augmentation methods in NER.
+[^6] As shown in Appendix I, not only the uncertainty performance but also the prediction performance could be affected by preserving the structure of a sentence.
 MELM has the second lowest perplexity after MR, and may be adversely affected by generated entities that are adapted to the context but not actually present.
@@
 While this experiment provided valuable insights into the impact of data augmentation on confidence calibration and uncertainty estimation in NER across different genres and languages, there are several limitations that should be acknowledged.
-First, due to resource limitations, the experiment was limited to evaluation with English as the source language.
+Source Language
+Due to resource limitations, the experiment was limited to evaluation with English as the source language.
 To effectively investigate the calibration and uncertainty of zero-shot cross-lingual transfer, it is important to expand the investigation to include a wider range of languages as the source language.
 Therefore, future research should prioritize the investigation of calibration and uncertainty performance using different languages as the source for zero-shot cross-lingual transfer.
-Second, as mentioned in Section 5.3, regarding the calibration and uncertainty evaluation policy, we simply evaluated an entity span as a single data instance, but a rigorous evaluation method that performs evaluation while considering multiple span candidates has been proposed Jiang et al.
+Evaluation of Uncertainty for Entities
+As mentioned in Section 5.3, regarding the calibration and uncertainty evaluation policy, we simply evaluated an entity span as a single data instance, but a rigorous evaluation method that performs evaluation while considering multiple span candidates has been proposed Jiang et al.
 (2022).
 Establishing span-level NER calibration evaluation methods that can efficiently and comprehensively evaluate calibration and uncertainty for entity types for datasets with many entity types and long sequence lengths is a topic for future research.
-Lastly, we broadly evaluated the calibration and uncertainty performance in both cross-genre and cross-lingual settings on data augmentation for NER, but only using sequence labeling-based methods.
-Recently, other paradigms in NER, such as span-based methods Fu et al.
-(2021) and Seq2Seq (sequence-to-sequence)-based methods Yan et al.
-(2021), have been proposed.
+NER Paradigm
+We broadly evaluated the calibration and uncertainty performance in both cross-genre and cross-lingual settings on data augmentation for NER, but only using sequence labeling-based methods.
+Recently, other paradigms in NER have been proposed such as the span-based methods Fu et al.
+(2021) and the generation-based methods Yan et al.
+(2021) including BART Lewis et al.
+(2020) or Large Language Models (LLM)  Xu et al.
+(2024), which are also applicable to nested-NER.
 In the future, the calibration or uncertainty performance of these methods could be evaluated.
+Other Data Augmentation Methods
+In this study, we focused on the data augmentation methods based on token or entity replacement.
+On the other hand, paraphrase-based data augmentation methods using such as LLM have attracted attention Ding et al.
+(2024).
+By using LLM, it is also possible to generate entities that correspond to a specified entity type Ye et al.
+(2024).
+To Investigate these in the context of uncertainty estimation also will be an interesting research.
 Ethical Considerations
@@
 The authors also acknowledge the Nara Institute of Science and Technology’s HPC resources made available for conducting the research reported in this paper.
-Appendix A Inference Time
-Methods Inference time [s]
-Baseline 14.90 $\pm$ 0.10
-TS 15.53 $\pm$ 0.92
-LS 14.94 $\pm$ 0.24
-MC Dropout 271.77 $\pm$ 1.81
-LwTR 14.91 $\pm$ 0.10
-MR 14.93 $\pm$ 0.17
-SR 14.83 $\pm$ 0.12
-MELM 14.89 $\pm$ 0.14
-Table 7: Inference time for each algorithm on MultiCoNER $\mathtt{EN}$ full test data.
-Methods OntoNotes 5.0 ($\mathtt{bc}$) OntoNotes 5.0 ($\mathtt{bn}$)
-$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$ $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
-Baseline 81.39$\pm$0.78 80.86$\pm$1.03 81.61$\pm$1.36 75.49$\pm$0.90 68.83$\pm$1.27 45.74$\pm$0.74 80.74$\pm$1.21 90.25$\pm$0.36 81.47$\pm$0.96 81.04$\pm$0.64 72.36$\pm$1.88 46.86$\pm$0.52
-TS 81.10$\pm$0.94 81.19$\pm$0.89 80.80$\pm$1.37 75.14$\pm$1.60 69.20$\pm$2.73 45.58$\pm$1.02 81.31$\pm$1.18 90.37$\pm$0.49 80.96$\pm$1.32 81.13$\pm$0.62 71.83$\pm$1.76 46.50$\pm$0.69
-LS 81.21$\pm$1.11 81.17$\pm$0.91 81.43$\pm$1.33 75.30$\pm$1.26 69.64$\pm$1.45 45.75$\pm$0.82 82.08$\pm$0.62 90.32$\pm$0.36 81.22$\pm$0.52 80.95$\pm$0.37 72.45$\pm$1.38 46.69$\pm$0.60
-MC Dropout 81.49$\pm$0.80 81.06$\pm$0.71 81.12$\pm$0.63 75.24$\pm$1.02 69.53$\pm$1.78 45.73$\pm$0.46 81.55$\pm$0.63 90.21$\pm$0.36 80.80$\pm$1.10 81.11$\pm$0.46 73.13$\pm$1.97 46.71$\pm$0.60
-LwTR (DA) 80.85$\pm$0.82 80.91$\pm$0.93 81.45$\pm$1.08 75.33$\pm$0.82 68.40$\pm$0.94 45.53$\pm$0.84 79.43$\pm$1.13 89.98$\pm$0.40 80.75$\pm$0.67 80.33$\pm$0.31 69.62$\pm$1.80 46.23$\pm$0.54
-MR (DA) 80.93$\pm$0.61 80.88$\pm$0.61 82.02$\pm$0.66 75.66$\pm$0.79 69.49$\pm$1.78 45.38$\pm$0.72 79.93$\pm$1.43 90.07$\pm$0.23 81.70$\pm$0.61 80.54$\pm$0.50 72.44$\pm$1.46 46.45$\pm$0.47
-SR (DA) 81.52$\pm$0.69 81.20$\pm$0.78 79.93$\pm$0.95 75.08$\pm$0.89 69.86$\pm$1.30 46.04$\pm$0.57 80.24$\pm$1.44 90.05$\pm$0.21 80.92$\pm$0.93 80.84$\pm$0.42 70.80$\pm$1.66 46.98$\pm$0.61
-MELM (DA) 81.08$\pm$0.37 80.81$\pm$0.97 80.11$\pm$0.98 74.74$\pm$1.24 66.68$\pm$1.18 45.19$\pm$1.05 79.23$\pm$0.64 90.26$\pm$0.38 81.48$\pm$0.65 80.66$\pm$0.79 68.42$\pm$1.65 46.36$\pm$0.44
-Methods OntoNotes 5.0 ($\mathtt{nw}$) OntoNotes 5.0 ($\mathtt{tc}$)
-$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$ $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
-Baseline 74.34$\pm$4.10 83.08$\pm$1.19 73.56$\pm$3.31 90.08$\pm$0.31 72.59$\pm$1.34 46.47$\pm$0.59 55.29$\pm$2.01 59.13$\pm$2.80 50.68$\pm$3.51 46.14$\pm$4.31 69.52$\pm$1.45 40.85$\pm$1.36
-TS 75.34$\pm$1.67 83.02$\pm$0.98 75.01$\pm$2.21 90.04$\pm$0.24 71.98$\pm$1.17 46.29$\pm$0.87 56.81$\pm$2.05 59.04$\pm$2.95 52.98$\pm$3.34 48.85$\pm$3.26 67.45$\pm$2.30 41.12$\pm$1.27
-LS 76.60$\pm$1.65 83.27$\pm$1.49 75.79$\pm$2.00 90.20$\pm$0.26 71.91$\pm$2.67 46.68$\pm$0.69 53.98$\pm$3.40 56.12$\pm$6.02 51.17$\pm$5.94 48.62$\pm$4.82 66.01$\pm$3.26 40.63$\pm$1.83
-MC Dropout 75.07$\pm$2.84 82.69$\pm$2.11 73.79$\pm$2.23 89.98$\pm$0.56 71.96$\pm$1.43 46.25$\pm$0.92 55.16$\pm$1.70 58.95$\pm$2.87 51.11$\pm$3.75 47.31$\pm$4.48 69.15$\pm$3.05 40.57$\pm$1.44
-LwTR (DA) 74.80$\pm$1.57 83.01$\pm$0.41 75.01$\pm$3.35 89.79$\pm$0.28 70.85$\pm$1.13 46.78$\pm$0.54 54.01$\pm$2.14 60.86$\pm$2.89 53.89$\pm$3.76 50.20$\pm$3.77 69.53$\pm$1.60 40.80$\pm$0.97
-MR (DA) 73.57$\pm$1.09 81.52$\pm$2.09 71.43$\pm$3.80 89.90$\pm$0.34 68.31$\pm$3.52 44.88$\pm$1.38 53.73$\pm$2.35 57.46$\pm$3.70 52.74$\pm$3.27 46.90$\pm$4.87 68.57$\pm$2.71 40.50$\pm$1.79
-SR (DA) 73.64$\pm$3.45 82.03$\pm$2.14 72.25$\pm$4.88 90.24$\pm$0.11 66.18$\pm$4.59 46.38$\pm$1.45 53.41$\pm$2.46 58.54$\pm$3.20 53.08$\pm$4.85 46.48$\pm$7.08 68.13$\pm$1.41 41.20$\pm$1.23
-MELM (DA) 73.46$\pm$2.46 82.22$\pm$1.23 75.56$\pm$2.60 89.94$\pm$0.18 62.43$\pm$2.95 45.19$\pm$0.97 48.01$\pm$5.27 49.59$\pm$6.16 48.93$\pm$4.11 42.09$\pm$5.61 63.46$\pm$2.28 36.16$\pm$3.76
-Table 8: F1 scores of existing calibration methods and data augmentation methods in OntoNotes 5.0.
+Appendix A Licenses of Datasets
+OntoNotes 5.0 can be used for research purposes as described in https://catalog.ldc.upenn.edu/LDC2013T19.
+MultiCoNER dataset is licensed by CC BY 4.0 as described in https://aws.amazon.com/marketplace/pp/prodview-cdhrtt7vq4hf4.
 Appendix B Details of Existing Calibration Methods
@@
 $y^{LS}_{i}=y_{i}(1-\epsilon)+\frac{\epsilon}{K}.$
+Methods Inference time [s]
+Baseline 14.90 $\pm$ 0.10
+TS 15.53 $\pm$ 0.92
+LS 14.94 $\pm$ 0.24
+MC Dropout 271.77 $\pm$ 1.81
+LwTR 14.91 $\pm$ 0.10
+MR 14.93 $\pm$ 0.17
+SR 14.83 $\pm$ 0.12
+MELM 14.89 $\pm$ 0.14
+Table 8: Inference time for each algorithm on MultiCoNER $\mathtt{EN}$ full test data.
 B.3 Monte-Carlo Dropout (MC Dropout)
@@
 $p_{i}=\frac{1}{M}\sum_{t=1}^{M}\frac{\exp(\bm{z}_{i}^{(t)})}{\sum_{j=1}^{K}\exp(\bm{z}_{j}^{(t)})}.$
-Methods $\mathtt{EN}$ $\mathtt{DE}$ $\mathtt{ES}$ $\mathtt{HI}$
-Baseline 68.80$\pm$0.38 64.91$\pm$0.60 63.53$\pm$0.41 37.33$\pm$3.77
-TS 68.51$\pm$0.52 64.70$\pm$0.90 63.41$\pm$0.45 37.90$\pm$2.79
-LS 69.17$\pm$0.55 65.37$\pm$0.51 63.83$\pm$0.32 39.93$\pm$3.50
-MC Dropout 68.56$\pm$0.96 64.70$\pm$0.87 63.39$\pm$0.69 36.38$\pm$5.89
-LwTR (DA) 68.86$\pm$0.82 64.95$\pm$0.64 63.52$\pm$0.85 38.24$\pm$3.11
-MR (DA) 69.71$\pm$0.72 65.37$\pm$0.57 64.25$\pm$0.62 37.53$\pm$4.03
-SR (DA) 68.81$\pm$0.41 64.75$\pm$0.86 63.85$\pm$0.46 42.31$\pm$1.45
-MELM (DA) 68.57$\pm$0.54 63.40$\pm$0.49 62.76$\pm$0.64 37.78$\pm$3.16
-Table 9: F1 scores of existing calibration methods and data augmentation methods in MultiCoNER.
-Appendix C F1 Scores
-Table 8 and 9 show F1 scores.
-Note that in many cases, data augmentation methods do not degrade predictive performance itself, but MELM often significantly degrades predictive performance in some cases, especially when the source domains are $\mathtt{nw}$ and $\mathtt{tc}$.
-Considering Section 6.1 and 6.2, MR improves calibration and uncertainty performance in many cases without degrading predictive performance.
-Appendix D More Results about Test Set Duplication
+Methods $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{nw}$ $\mathtt{tc}$
+ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓)
+Baseline 27.07 33.52 26.08 31.17 26.66 31.35 37.66 46.32
+TS 27.25 33.41 26.17 31.17 26.68 31.34 36.66 45.52
+LS 27.19 33.57 25.88 30.49 26.52 30.67 35.24 43.68
+MC Dropout 27.15 33.61 25.90 30.85 26.62 31.18 36.80 45.71
+LwTR (DA) 27.65 33.78 26.49 31.78 27.28 31.67 35.90 44.97
+MR (DA) 27.33 33.22 26.21 31.00 26.26 30.53 36.38 44.65
+SR (DA) 27.23 33.08 26.11 30.72 27.47 31.89 35.24 43.57
+MELM (DA) 27.95 33.88 26.63 30.91 27.62 32.09 34.83 42.65
+Table 9: ECE and MCE averaged over all target domain results in OntoNotes 5.0.
+Appendix C Inference Time
+Table 8 shows the results of the inference time on MultiCoNER $\mathtt{EN}$ set. We can see that data augmentation methods do not affect the computational overhead during inference clearly.
+Appendix D Full Averaged Results on OntoNotes 5.0
+To briefly summarize the many values in Table 3 and  4, we averaged the ECE and MCE scores for each method and domain, shown them in Table 9.
+From this table, we can see that data augmentation methods are slightly worse than existing methods some cases when averaging all settings , while in others, especially $\mathtt{nw}$ and $\mathtt{tc}$, data augmentation methods are better on average.
+Appendix E More Results about Test Set Duplication
 Table 10 shows the results of the percentage increase in entity duplication that are new overlaps with each target domain’s test set when applying each data augmentation method except MR, where the source domains are $\mathtt{bc}$, $\mathtt{bn}$, and $\mathtt{nw}$.
@@
 Table 12: The percentage increase in entity duplication in the case of the source domain $\mathtt{tc}$ that are new overlaps with each target domain’s test set when applying each data augmentation method except MR.
-More results are in Appendix D.
-Appendix E Impact of New Entities via Data Augmentation
+More results are in Appendix E.
+Appendix F Impact of New Entities via Data Augmentation
 To investigate the impact of new entities added by data augmentation methods on calibration performance, we measured the percentage of new entities added in the training data and the percentage of new entities that overlap with the test set.
@@
 Therefore, new entities by data augmentation methods for NER are likely to have no effect on calibration performance or uncertainty performance.
-Appendix F t-SNE Plot for MultiCoNER Dataset
+Appendix G t-SNE Plot for MultiCoNER Dataset
 To overview of the ID and OOD data instances in the MultiCoNER dataset, t-SNE plot is shown in Figure 4.
-Appendix G Results for Low-resource Language
+Appendix H Results for Low-resource Language
 To investigate the uncertainty estimation performance for low-resource language, we additionaly show the results of 10,000 examples of Bangla ($\mathtt{BN}$) from MultiCoNER dataset in Table 13 when source language is $\mathtt{EN}$.
@@
 Table 13: Results of existing methods and data augmentation methods in MultiCoNER $\mathtt{BN}$.
-Appendix H Licenses of Datasets
-OntoNotes 5.0 can be used for research purposes as described in https://catalog.ldc.upenn.edu/LDC2013T19.
-MultiCoNER dataset is licensed by CC BY 4.0 as described in https://aws.amazon.com/marketplace/pp/prodview-cdhrtt7vq4hf4.
+Methods OntoNotes 5.0 ($\mathtt{bc}$) OntoNotes 5.0 ($\mathtt{bn}$)
+$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$ $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
+Baseline 81.39$\pm$0.78 80.86$\pm$1.03 81.61$\pm$1.36 75.49$\pm$0.90 68.83$\pm$1.27 45.74$\pm$0.74 80.74$\pm$1.21 90.25$\pm$0.36 81.47$\pm$0.96 81.04$\pm$0.64 72.36$\pm$1.88 46.86$\pm$0.52
+TS 81.10$\pm$0.94 81.19$\pm$0.89 80.80$\pm$1.37 75.14$\pm$1.60 69.20$\pm$2.73 45.58$\pm$1.02 81.31$\pm$1.18 90.37$\pm$0.49 80.96$\pm$1.32 81.13$\pm$0.62 71.83$\pm$1.76 46.50$\pm$0.69
+LS 81.21$\pm$1.11 81.17$\pm$0.91 81.43$\pm$1.33 75.30$\pm$1.26 69.64$\pm$1.45 45.75$\pm$0.82 82.08$\pm$0.62 90.32$\pm$0.36 81.22$\pm$0.52 80.95$\pm$0.37 72.45$\pm$1.38 46.69$\pm$0.60
+MC Dropout 81.49$\pm$0.80 81.06$\pm$0.71 81.12$\pm$0.63 75.24$\pm$1.02 69.53$\pm$1.78 45.73$\pm$0.46 81.55$\pm$0.63 90.21$\pm$0.36 80.80$\pm$1.10 81.11$\pm$0.46 73.13$\pm$1.97 46.71$\pm$0.60
+LwTR (DA) 80.85$\pm$0.82 80.91$\pm$0.93 81.45$\pm$1.08 75.33$\pm$0.82 68.40$\pm$0.94 45.53$\pm$0.84 79.43$\pm$1.13 89.98$\pm$0.40 80.75$\pm$0.67 80.33$\pm$0.31 69.62$\pm$1.80 46.23$\pm$0.54
+MR (DA) 80.93$\pm$0.61 80.88$\pm$0.61 82.02$\pm$0.66 75.66$\pm$0.79 69.49$\pm$1.78 45.38$\pm$0.72 79.93$\pm$1.43 90.07$\pm$0.23 81.70$\pm$0.61 80.54$\pm$0.50 72.44$\pm$1.46 46.45$\pm$0.47
+SR (DA) 81.52$\pm$0.69 81.20$\pm$0.78 79.93$\pm$0.95 75.08$\pm$0.89 69.86$\pm$1.30 46.04$\pm$0.57 80.24$\pm$1.44 90.05$\pm$0.21 80.92$\pm$0.93 80.84$\pm$0.42 70.80$\pm$1.66 46.98$\pm$0.61
+MELM (DA) 81.08$\pm$0.37 80.81$\pm$0.97 80.11$\pm$0.98 74.74$\pm$1.24 66.68$\pm$1.18 45.19$\pm$1.05 79.23$\pm$0.64 90.26$\pm$0.38 81.48$\pm$0.65 80.66$\pm$0.79 68.42$\pm$1.65 46.36$\pm$0.44
+Methods OntoNotes 5.0 ($\mathtt{nw}$) OntoNotes 5.0 ($\mathtt{tc}$)
+$\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$ $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{mz}$ $\mathtt{nw}$ $\mathtt{tc}$ $\mathtt{wb}$
+Baseline 74.34$\pm$4.10 83.08$\pm$1.19 73.56$\pm$3.31 90.08$\pm$0.31 72.59$\pm$1.34 46.47$\pm$0.59 55.29$\pm$2.01 59.13$\pm$2.80 50.68$\pm$3.51 46.14$\pm$4.31 69.52$\pm$1.45 40.85$\pm$1.36
+TS 75.34$\pm$1.67 83.02$\pm$0.98 75.01$\pm$2.21 90.04$\pm$0.24 71.98$\pm$1.17 46.29$\pm$0.87 56.81$\pm$2.05 59.04$\pm$2.95 52.98$\pm$3.34 48.85$\pm$3.26 67.45$\pm$2.30 41.12$\pm$1.27
+LS 76.60$\pm$1.65 83.27$\pm$1.49 75.79$\pm$2.00 90.20$\pm$0.26 71.91$\pm$2.67 46.68$\pm$0.69 53.98$\pm$3.40 56.12$\pm$6.02 51.17$\pm$5.94 48.62$\pm$4.82 66.01$\pm$3.26 40.63$\pm$1.83
+MC Dropout 75.07$\pm$2.84 82.69$\pm$2.11 73.79$\pm$2.23 89.98$\pm$0.56 71.96$\pm$1.43 46.25$\pm$0.92 55.16$\pm$1.70 58.95$\pm$2.87 51.11$\pm$3.75 47.31$\pm$4.48 69.15$\pm$3.05 40.57$\pm$1.44
+LwTR (DA) 74.80$\pm$1.57 83.01$\pm$0.41 75.01$\pm$3.35 89.79$\pm$0.28 70.85$\pm$1.13 46.78$\pm$0.54 54.01$\pm$2.14 60.86$\pm$2.89 53.89$\pm$3.76 50.20$\pm$3.77 69.53$\pm$1.60 40.80$\pm$0.97
+MR (DA) 73.57$\pm$1.09 81.52$\pm$2.09 71.43$\pm$3.80 89.90$\pm$0.34 68.31$\pm$3.52 44.88$\pm$1.38 53.73$\pm$2.35 57.46$\pm$3.70 52.74$\pm$3.27 46.90$\pm$4.87 68.57$\pm$2.71 40.50$\pm$1.79
+SR (DA) 73.64$\pm$3.45 82.03$\pm$2.14 72.25$\pm$4.88 90.24$\pm$0.11 66.18$\pm$4.59 46.38$\pm$1.45 53.41$\pm$2.46 58.54$\pm$3.20 53.08$\pm$4.85 46.48$\pm$7.08 68.13$\pm$1.41 41.20$\pm$1.23
+MELM (DA) 73.46$\pm$2.46 82.22$\pm$1.23 75.56$\pm$2.60 89.94$\pm$0.18 62.43$\pm$2.95 45.19$\pm$0.97 48.01$\pm$5.27 49.59$\pm$6.16 48.93$\pm$4.11 42.09$\pm$5.61 63.46$\pm$2.28 36.16$\pm$3.76
+Table 14: $F_{1}$ scores of existing calibration methods and data augmentation methods in OntoNotes 5.0.
+Methods $\mathtt{EN}$ $\mathtt{DE}$ $\mathtt{ES}$ $\mathtt{HI}$
+Baseline 68.80$\pm$0.38 64.91$\pm$0.60 63.53$\pm$0.41 37.33$\pm$3.77
+TS 68.51$\pm$0.52 64.70$\pm$0.90 63.41$\pm$0.45 37.90$\pm$2.79
+LS 69.17$\pm$0.55 65.37$\pm$0.51 63.83$\pm$0.32 39.93$\pm$3.50
+MC Dropout 68.56$\pm$0.96 64.70$\pm$0.87 63.39$\pm$0.69 36.38$\pm$5.89
+LwTR (DA) 68.86$\pm$0.82 64.95$\pm$0.64 63.52$\pm$0.85 38.24$\pm$3.11
+MR (DA) 69.71$\pm$0.72 65.37$\pm$0.57 64.25$\pm$0.62 37.53$\pm$4.03
+SR (DA) 68.81$\pm$0.41 64.75$\pm$0.86 63.85$\pm$0.46 42.31$\pm$1.45
+MELM (DA) 68.57$\pm$0.54 63.40$\pm$0.49 62.76$\pm$0.64 37.78$\pm$3.16
+Table 15: $F_{1}$ scores of existing calibration methods and data augmentation methods in MultiCoNER.
+Appendix I $F_{1}$ Scores
+Table 14 and 15 show $F_{1}$ scores.
+Note that in many cases, data augmentation methods do not degrade predictive performance itself, but MELM often significantly degrades predictive performance in some cases, especially when the source domains are $\mathtt{nw}$ and $\mathtt{tc}$.
+Considering Section 6.1 and 6.2, MR improves calibration and uncertainty performance in many cases without degrading predictive performance.